XYZ Company’s machine was purchased 5 years ago for $55,000. It had an expected life of 10 years when it was bought, and its remaining depreciation is $5,500 per year for each year of its remaining life and can be sold for $20,000 at the end of its useful life. A new machine can be purchased for $120,000, including the installation costs. During its 5-year life, it will reduce cash operating expenses by $30,000 per year. Sales revenue will not be affected. At the end of its useful life, the machine is estimated to be sold at $10,000. We will use MARCS depreciation, and the machine will be depreciated over its 5-year property class life. The old machine can be sold today for $35,000.

*The tax rate is 35%

*WACC is 16%

a) what is the amount of the initial cash flow at Year 0 if the new machine is purchased?

b)Calculate the after-tax salvage value of the new machine at the end of the project?

c)Calculate the incremental cash flows that will occur at the end of years 1-5?

*Use excel cell reference for the questions for the above questions
Answer #1
a) Cost of the new machine $   1,20,000
Sale value of the old machine $      35,000
Book value = 5500*5 = $      27,500
Gain on sale $        7,500
Tax on gain at 35% $        2,625
After tax sale value of old machine = 35000-2625 = $       32,375
Initial cash flow at Year 0 $       87,625
b) 0 1 2 3 4 5
Savings in operating expenses $       30,000 $     30,000 $      30,000 $     30,000 $      30,000
Incremental depreciation:
Depreciation on new machine [MACRS 5 Year] $       24,000 $     38,400 $      23,040 $     13,824 $      13,824 6912
Depreciation on the old machine $         5,500 $       5,500 $         5,500 $        5,500 $        5,500
Incremental depreciation $       18,500 $     32,900 $      17,540 $        8,324 $        8,324
Incremental NOI $       11,500 $      -2,900 $      12,460 $     21,676 $      21,676
Tax at 35% $         4,025 $      -1,015 $         4,361 $        7,587 $        7,587
Incremental NOPAT $         7,475 $      -1,885 $         8,099 $     14,089 $      14,089
Add: Incremental depreciation $       18,500 $     32,900 $      17,540 $        8,324 $        8,324
Incremental OCF $       25,975 $     31,015 $      25,639 $     22,413 $      22,413
Incremental salvage value:
After tax salvage value of new machine = 10000-(10000-6912)*35% = $        8,919
After tax salvage value of old machine = 20000*(1-35%) = $      13,000
Incremental after tax salvage value = 8919-13000 = $      -4,081
Incremental project cash flows $       25,975 $     31,015 $      25,639 $     22,413 $      18,333
d) PVIF at 16% {PVIF = 1/1.16^t] 0.86207 0.74316 0.64066 0.55229 0.47611
PV at 16% $       22,392 $     23,049 $      16,426 $     12,379 $        8,728
Total PV of CFS t1 to t5 $      82,974
Less: Initial investment $      87,625
NPV $       -4,651
e) As the NPV is negative, the replacement should not be done.